# Choose Your Model Size: Any Compression by a Single Gradient Descent

### v1

@@ 4 Experiments
To demonstrate effectiveness across architectural differences in LLMs, we evaluate ACIP on a range of popular open-weight models: LLaMA-7B/13B (Touvron et al., 2023a), LLaMA-2-7B/13B (Touvron et al., 2023b), LLaMA-3.1-8B (Grattafiori et al., 2024), Qwen2.5-7B/14B (Qwen et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023). More detailed analyses and ablations are carried out with LLaMA-7B as it was most extensively studied in previous research on structured weight pruning. Regarding evaluation tasks, we follow (Wang et al., 2024) and report perplexity on validation held-outs of C4 (Raffel et al., 2019) and WikiText-2 (Merity et al., 2017), and consider seven zero-shot tasks from EleutherAI LM Evaluation Harness (LM-Eval) (Gao et al., 2023). Implementation details about ACIP and choices of hyperparameters can be found in Appendix B.
4.1 Analyzing Compression-Performance Trade-Offs
Picking up the introductory discussion of Figure 1, we now study compression-performance trade-offs of ACIP in more detail. Figure 4 and Figure 5 confirm smooth and consistent curve shapes for all considered models; analogous results for Wikitext-2 and individual zero-shot LM-Eval tasks can be found in Figure A8. A remarkable observation is that the oldest models, LLaMA-7B/13B, perform best perplexity-wise, while newer, more capable models like Qwen2.5-7B/14B dominate on LM-Eval as expected, especially on the lower compression levels. This apparent contradiction is likely caused by a deviation of the LLM pre-training data distributions from C4. In fact, more recent models use much larger and more diverse training corpora.
A second noteworthy outcome of Figure 4 and Figure 5 are the gaps between LLMs of different base model sizes in the same family. Indeed, ACIP cannot match the performance of base models of smaller size, e.g., compare the compressed Qwen2.5-7B with the original Qwen2.5-3B. This is not surprising because the corresponding smaller-size base models were obtained by pre-training or knowledge distillation, which are orders of magnitudes more expensive than ACIP.
Comparison to Existing Works
We now compare ACIP to two recent works focusing on SVD-based structured pruning, namely ASVD (Yuan et al., 2024) and SVD-LLM (Wang et al., 2024). Both approaches are backpropagation-free and perform (activation-aware) layer-wise updates instead. Table 1 shows that ACIP consistently outperforms both methods with a growing gap for higher compression levels. Note that SVD-LLM was calibrated on Wikitext-2 instead of C4, which might explain slightly better results on the former dataset for 70% and 80% size. We think that these results underpin the benefits of an end-to-end scheme: (i) a simultaneous correction, e.g., by LoRA, can drastically improve performance, and (ii) robust pruning patterns can be found without leveraging any specific features of the SVD factorization. Moreover, note that recomputations are required to generate each row of Table 1 for ASVD and SVD-LLM, whereas ACIP only needs a single run. Analogous results for ACIP applied to all other models can be found in Table A2.
@@ 4.2 Improving Performance Through Fine-Tuning
While the main goal of this work is to produce a full family of accurate, compressed models from a few optimization steps, their performance can be certainly improved through continued fine-tuning. Figure 6 highlights the gains of fine-tuning LLaMA-7B; see Table A2 for more detailed numerical results on all other models. We observe that fine-tuning leads to a performance offset that is almost constant across all compression levels, which underlines the predictive capacity of ACIP. Note that we even observe a jump at zero compression because inserting the low-rank adapter learned by ACIP leads to a slight initial performance drop.
We argue that post-compression fine-tuning should be still seen as an independent (and much more costly) algorithmic step for two reasons. (i) Its outcome strongly depends on the specific training protocol and data, making a fair comparison challenging (cf. Q3 in Appendix A); (ii) it requires us to fix a compression level, which breaks the crucial any-compression-feature of ACIP.
@@ 4.3 Combining ACIP with Quantization
In the field of low-cost compression for LLMs, quantization is still considered as the gold standard (Hohman et al., 2024; Zhu et al., 2024), so that a practitioner might not be willing to exchange its gains for the benefits of ACIP. Fortunately, ACIP only tunes a tiny fraction of weights with high precision, so that all remaining modules are suitable for quantization. In our experiments, we quantize all parameterized and unparametrized linear layers to $4$-bit in fp4-format (Dettmers et al., 2023) using the bitsandbytes-Package (W4A16), except for the embedding layer and final classification head. We study the gains of quantization for ACIP in the following two ways.
@@
We first apply ACIP as usual, compress the model to a given target size, and then quantize all remaining linear layers. Figure 6 confirms that this approach works fairly well, only producing a slight performance drop compared to non-quantized versions; see Table A3 for a full evaluation on all other metrics. We also observe that an optional fine-tuning step as in Section 4.2 can almost fully compensate for the errors introduced by quantization after compression. This finding is well in line with the effectiveness of the popular QLoRA approach (Dettmers et al., 2023). Moreover, Figure 6 reveals a drastic improvement through quantization in terms of required memory. Here, the ACIP-trade-off allows practitioners to study and apply a more fine-grained compression on top of quantization.
@@
Compared to layer-wise methods like ASVD and SVD-LLM, ACIP has a higher demand in GPU memory due to backpropagation. A quantization of all frozen weight matrices can be an effective remedy in this respect. For the experiment shown in Figure 7, we have applied quantization before ACIP, which leads to very similar compression-performance trade-offs as in the non-quantized case. Going one step further, we transfer the score maps and low-rank adapters from this quantized version of ACIP back to full precision: We load the base model in bf16, apply layer-wise SVD-parametrization, insert the low-rank adapters learned by quantized ACIP, and use the corresponding score map to obtain a compressed model (W16A16). The resulting trade-off curve in Figure 7 confirms that this simple strategy works fairly well, especially for lower compression levels.
Further Experiments and Ablations
Several additional experiments are presented in Appendix D, analyzing the impact of several key components and design aspects of ACIP.
@@ 5 Conclusion
In this work, we have introduced Any Compression via Iterative Pruning (ACIP), a simple end-to-end algorithm to determine the compression-performance trade-off of large pre-trained models. The underlying score map ranking allows us to materialize models of any compression rate from a single optimization run. We have demonstrated empirically that the downstream performance of the resulting models is superior to existing, layer-wise factorization approaches. The flexibility and efficiency of ACIP make it a practical tool for deploying large-scale models in resource-constrained settings, especially in combination with other compression techniques such as quantization.
Limitations and Outlook
Our main results in Figure 4 and Figure 5 resemble the well-known phenomenon of scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). We believe that this association is not a coincidence. But revealing a more rigorous connection goes beyond the scope of this work. In a similar vein, we observe that more recent models tend to be less compressible (e.g., compare the slopes of LLaMA-13B and Qwen2.5-14B in Figure 5). We hypothesize that this relates to the fact that newer models carry more dense information per weight, since they were trained on much larger datasets (Allen-Zhu & Li, 2024). Also, the distribution of the calibration dataset (C4 in our case) might play an important role in this context.
A notable limitation of our work is that we have only focused on models that are tunable on a single (NVIDIA H100) GPU in bf16-precision. Hence, the scaling behavior of ACIP for larger LLMs (30B+) remains to be explored. Finally, we emphasize that ACIP could be transferred to other modalities, architectures, and tasks without any notable modifications.
@@
Code will be made available on Github soon.
Acknowledgements
@@
We kindly acknowledge funding by the German Federal Ministry of Education and Research (BMBF) within the project “More-with-Less: Effiziente Sprachmodelle für KMUs” (grant no. 01IS23013A). Moreover, we thank WestAI for providing compute resources that were used to conduct the numerical experiments of this work.
Appendix A Question & Answers
In this section, we discuss a few common questions about our method and experimental design that were not (fully) addressed in the main part due to length restrictions.
@@ Appendix A Question & Answers
    We argue that these are fundamentally different compression approaches. Full weight manipulations, in principle, have the potential to lead to more powerful compressions because they have more degrees of freedom (analogously to full-weight fine-tuning vs. PEFT). Therefore, they should not be seen as competing methods but complementary ones. We admit that practitioners probably would not favor ACIP over well-established and widely supported quantization techniques. However, the adapter-style nature of ACIP makes it suitable for a combination which can lead to extra gains, as demonstrated in Section 4.3.
@@
    Why did you not compare with model distillation or combined ACIP with it?
@@
    While model distillation can lead to outstanding compression results, e.g., see (Raschka, 2024), this approach requires significantly more resources than ACIP, typically orders of magnitudes more. A direct comparison is therefore not meaningful from our point of view, as it should at least be based on approximately the same computational budget.
@@
    Why did you not compare your fine-tuning results with other factorization-based methods?
@@
    Some previous works on SVD-based pruning (Wang et al., 2024; Jaiswal et al., 2024) report results on post-compression fine-tuning as well. While our own results in Table A2 are highly competitive in this respect, we decided to omit a direct comparison because: (i) the downstream performance strongly depends on the exact fine-tuning protocol (datasets, compute budget, hyperparameters, etc.) which we were not able to fully recover from published code; and (ii) as pointed out in Section 4.2, promoting a costly fine-tuning step after compression is not the primary concern of our work.
@@
    Let us first summarize several benefits of our end-to-end optimization approach from the main paper: (i) it is conceptually simple and requires no feature engineering, (ii) an error correction can be injected with almost no extra costs, (iii) it allows us to perform efficient and accurate any-size compression.
    Apart from that, and to the best of our knowledge, existing compression algorithms that use layer-wise updates like ASVD (Yuan et al., 2024), SVD-LLM (Wang et al., 2024), or WeLore (Jaiswal et al., 2024) require a separate fine-tuning step to achieve competitive downstream performance. Therefore, the lower costs of layer-wise compression are actually dominated by a more expensive backpropagation-based step. It remains open if similar results can be obtained by a fully tuning-free algorithm.
  • Q9.
    Why do you use matrix factorization, and SVD in particular?
  • A10.
    Committing to a backpropagation-based algorithm (see Q4) means that we have to deal with increased memory requirements. As such, matrix factorization is not helpful in that respect because the number of parameters might even increase initially (for instance, an SVD-parametrization basically doubles the size of a quadratic weight matrix). On the other hand, tuning and pruning only the bottleneck layer (i.e., the singular value mask in case of ACIP) has the potential for drastic size reductions and is highly parameter-efficient. For example, the number of tunable mask parameters for LLaMA-7B with ACIP is $<$1M.
    With this in mind, SVD as a specific matrix factorization is an obvious candidate due to its nice mathematical and numerical properties, in particular, optimal low-rank matrix approximation and stable matrix operations due to orthogonality.
@@ Appendix B Implementation Details
Following previous work on LLM compression, we use C4 for training as it is a good proxy of a general-purpose dataset. In the context of ACIP, it should be primarily seen as a calibration dataset that allows us to propagate meaningful activations through a pre-trained model while performing structured pruning. Overfitting to the distribution of C4 is implicitly mitigated, since we only tune very few parameters (masks and LoRA-adapters) compared to the total model size.
@@
All considered (evaluation) datasets and pre-trained models are imported with the HuggingFace transformers-library in bfloat16-precision. Our experiments were implemented with PyTorch (Paszke et al., 2019) and the Lightning package.
ACIP-Specifics
@@
As mentioned in Remark 2.1, we apply a linear scheduler that increases the regularization parameter $\lambda$ dynamically over the pruning process. This ensures that the pruning becomes more and more aggressive over time and the stopping criterion will be reached at some point. Across all experiments, we use $\lambda=1e{-}3$ as initial value and increase it by a factor of $1.01$ every $4$ steps (this amounts to a doubling $\lambda$ at about every $280$ steps).
@@
As pointed out in Section 2.2.2, we choose a (maximum) target compression rate as a stopping criterion for ACIP. In most experiments, a rate of $r_{\text{stop}}=0.4$ is reasonable (i.e., only 40% or the original parameters remain), and we refer to Section D.2 for further discussion and analysis. After the stopping criterion is reached with tune the low-rank adapter for 1K more steps while the masks are frozen (see Section 2.2.2).
@@
The mask parameters in (7) are rescaled by a fixed factor of $0.02$ to ensure a better alignment with the numerical range of the remaining network weights. The low-rank adapters are created with $r=32$, $\alpha=16$, and dropout $0.05$. For LLaMA-7B, the number of tunable parameters amounts to $<$1M mask parameters and approximately 80M low-rank adapter parameters.
@@
For sample data from C4, we use $1024$ tokens per sample and a batch size of $4$. We use Adam (Kingma & Ba, 2015) as optimizer without weight decay and a learning rate of $55e{-}$ 5.
Runtime Analysis
@@
ACIP requires significantly fewer steps than fine-tuning. Depending on when the stopping criterion is reached, it typically takes 1.5K - 2.5K steps, including 1K post-tuning steps of the low-rank adapters. For LLaMA-7B, for example, this amounts to a wall clock runtime of $<30$ minutes, including the initial SVD computations for the base model parametrization. All runs were performed on single NVIDIA H100 GPUs.
Fine-Tuning
@@
In all post-compression fine-tunings (see Section 4.2), we simply continue training ACIP’s low-rank adapters (the optimizer states are reset). We train for $20$K-$25$K steps on C4 with a batch size of $4$ and a learning rate of $42e{-}$ 4.
Appendix C Supplementary Results for Section 4.1Section 4.3
@@
Figure A8 complements the trade-off curves in Figure 4 and Figure 5 by all other considered evaluation metrics (see Section 4.1). Table A2 reports these results in terms of numbers, including all fine-tuning results for all models (see Section 4.2). Table A3 provides more detailed evaluation results on fine-tuning a quantized and compressed LLaMA-7B model (see Section 4.3).
@@
Table A2: Evaluation results for ACIP on all considered LLMs. Scores on C4 and WikiText-2 are measured in perplexity (smaller is better), and the LM-Eval zero-shot tasks are measured in accuracy (higher is better). The results of LLaMA2-13B were achieved by ignoring all up-projection layers in ACIP (see Section D.6 for more details).
@@
Original 9.99 5.05 0.56 0.82 0.63 0.53 0.35 0.81 0.75 0.64
(a) WikiText-2
(b) ARC Challenge
(c) ARC Easy
(d) HellaSwag
(e) MathQA
(f) OpenBookQA
(g) PIQA
(h) Winogrande
Figure A8: Compression-performance trade-off curves generated by ACIP on WikiText-2 and individual LM-Eval tasks, complementing the results of Figure 4 and Figure 5.
Table A3: More detailed evaluation results for our quantization experiments from Section 4.3, reported in terms of numbers.
Eff. model size [GB] C4 ↓ WikiText-2 ↓ ARC_c ↑ ARC_e ↑ HellaS. ↑ MathQA ↑ Openb. ↑ PIQA ↑ WinoG. ↑ LM Eval Avg. ↑
Size Ablation

### v2

@@ 3.3 Improving Performance Through Fine-Tuning
While the main goal of this work is to produce a full family of accurate, compressed models from a few optimization steps, their performance can be certainly improved through continued fine-tuning. Figure˜6 highlights the gains of fine-tuning LLaMA-7B; see Table˜A2 for more detailed numerical results on all other models. We observe that fine-tuning leads to a performance offset that is almost constant across all compression levels, which underlines the predictive capacity of ACIP. Note that we even observe a jump at zero compression because inserting the low-rank adapters learned by ACIP leads to a slight initial performance drop (see Section˜D.3 for a potential improvement).
An optional fine-tuning step is not exclusive to ACIP but can be applied to many other compression approaches as well. Table˜A3 provides a comparison with ASVD Yuan et al. (2024) and SVD-LLM Wang et al. (2024) (cf. Section˜3.2) when fine-tuned with LoRA. While ACIP still performs best in this respect, we argue that post-compression fine-tuning should be still seen as an independent (and much more costly) algorithmic step for two reasons. (i) Its outcome strongly depends on the specific training protocol and data, making a fair and direct comparison challenging; (ii) it requires us to fix a compression level, which breaks the crucial Any Compression feature of ACIP. Therefore, promoting a costly fine-tuning step after compression is not the primary concern of our work.
@@ 3.4 Combining ACIP with Quantization
In the field of low-cost compression for LLMs, quantization is still considered as the gold standard Hohman et al. (2024); Zhu et al. (2024), so that a practitioner might not be willing to exchange its gains for the benefits of ACIP. Fortunately, ACIP only tunes a tiny fraction of weights with high precision, so that all remaining modules are suitable for quantization. In our experiments, we quantize all parameterized and unparametrized linear layers to $4$-bit in fp4-format Dettmers et al. (2023) using the bitsandbytes-Package (W4A16), except for the embedding layer and final classification head. We study the gains of quantization for ACIP in the following two ways.
@@
We first apply ACIP as usual, compress the model to a given target size, and then quantize all linear layers. Figure˜6 confirms that this approach works fairly well, only producing a slight performance drop compared to non-quantized versions; see Table˜A4 for a full evaluation on all other metrics. We also observe that an optional fine-tuning step as in Section˜3.3 can almost fully compensate for the errors introduced by quantization after compression. This finding is well in line with the effectiveness of the popular QLoRA approach Dettmers et al. (2023). Moreover, Figure˜6 reveals a drastic improvement through quantization in terms of required memory. Here, the ACIP-trade-off allows practitioners to study and apply a more fine-grained compression on top of quantization.
@@
Compared to layer-wise methods like ASVD and SVD-LLM, ACIP has a higher demand in GPU memory due to backpropagation. A quantization of all frozen weight matrices can be an effective remedy in this respect. For the experiment shown in Figure˜7, we have applied quantization before ACIP, which leads to very similar compression-performance trade-offs as in the non-quantized case. Going one step further, we transfer the score maps and low-rank adapters from this quantized version of ACIP back to full precision: We load the base model in bf16, apply layer-wise SVD-parametrization, insert the low-rank adapters learned by quantized ACIP, and use the corresponding score map to obtain a compressed model (W16A16). The resulting trade-off curve in Figure˜7 confirms that this simple strategy works fairly well, especially for lower compression levels.
3.5 Further Experiments and Ablations
Several additional experiments are presented in Appendix˜D, analyzing the impact of several key components and design aspects of ACIP. Starting with an analysis of algorithmic efficiency and latency (Section˜D.1), we study the impact of the low-rank adapters (Section˜D.2), compression rule (Section˜D.3), stopping criterion (Sections˜D.4 and D.5), score map design (Section˜D.6), post-tuning (Section˜D.7), and specific types of linear layers (Section˜D.8). Finally, we show examples of score maps (Section˜D.9) and prompt completions by compressed models (Section˜D.10).
4 Related Work
In this section, we pick up on our broader discussion on the field of model compression from Section˜1 and put our work in context with several directly related branches of research.
Structured pruning & low-rank factorization.
Conceptually, ACIP falls under the umbrella of structured parameter pruning, specifically, low-rank matrix decomposition. The rationale behind this compression approach is to approximate large weight matrices by products of low-rank factors to reduce the total parameter count, and at the same time, to preserve critical information. After initial efforts into this direction for smaller language models Edalati et al. (2022); Tahaei et al. (2022), techniques for LLMs primarily built on (weighted) SVD of linear layers Ben Noach & Goldberg (2020); Hsu et al. (2022).
However, a key challenge of SVD-based pruning is that simply truncating singular values based on magnitude alone is insufficient and makes additional fine-tuning on downstream tasks necessary. Follow-up work recognized that the poor approximations are caused by LLM weights being high-rank and instead turned to decomposing network features which are sparse Kaushal et al. (2023); Yu & Wu (2023). Similarly, recent studies Sharma et al. (2023); Yuan et al. (2024); Jaiswal et al. (2024) have shown rank reduction to differently affect layers in a network and proposed heuristics for non-uniform pruning. Going even further, ASVD Yuan et al. (2024) pointed out the importance of activation-aware approximations, proposing a training-free compression method that takes the (calibration) data distribution into account. Building on this, SVD-LLM Wang et al. (2024) recently derived an analytical layer-wise correction leading to superior compression results. While ACIP promotes an activation-aware solution as well, it relies on gradient-based optimization, which avoids any SVD-specific feature engineering and allows for simultaneous errors corrections.
Another relevant line of work on structured pruning aims to jointly remove groups of parameters or entire network components, e.g., weight matrix columns/rows, network layers, or attention heads Frantar & Alistarh (2023); Ma et al. (2023); Xia et al. (2024); Ashkboos et al. (2024); Kim et al. (2024). At high compression rates, however, such “coarse” approaches often remove critical substructures, causing a significant performance drop that is only recoverable through additional fine-tuning.
Score maps & Any Compression.
A common feature of the aforementioned structured pruning approaches is that they first truncate parameters to a preset target size and then compute an error correction. This design choice means that exploring the full compression-performance trade-off requires repeated, costly computations for each desired compression ratio (cf. Figure˜1(a)). In contrast, ACIP overcomes this limitation by using score maps to determine global parameter importance. As such, score maps have been used as a tool for model compression since the 1980s LeCun et al. (1989); Hassibi et al. (1993). However, in the era of LLMs, deriving a score for each parameter poses significant challenges in terms of scalability. Addressing this concern, ACIP enables scalable Any Compression by (i) using weight factorization to significantly reduce the score map size (e.g., to $\sim$900k parameters for a 7B model), and (ii) decoupling the scoring and compression stages (Step 2 and 3 in Section˜2.2, respectively).
Any-size pre-training.
Beyond post-training compression, an alternative paradigm for obtaining models of varying sizes is to incorporate this flexibility into the (pre-)training process itself. For example, Cai et al. (2020) train a single, large “Once-for-all” network from which specialized sub-networks of different sizes can be extracted without retraining. More recently, the MatFormer Devvrit et al. (2024) has demonstrated how to build a family of “any-size” models directly through a nested pre-training methodology, based on Matryoshka Representation Learning Kusupati et al. (2022). The recently published Gemma-3n model uses this approach in production to make LLMs ready for mobile and edge devices Gonzalez & Shivanna (2025). While these methods also produce a trade-off between model size and performance, they require a significantly higher upfront computational budget associated with complex pre-training from scratch. ACIP provides a lightweight, post-training alternative that offers similar flexibility for any existing pre-trained model.
Rate-distortion theory.
Finally, our work can be viewed through the lens of rate-distortion theory, which investigates the analytical trade-off between achievable data compression rates and the error (distortion) introduced by lossy compression (Cover & Thomas, 2006). While some recent work (Gao et al., 2019; Isik et al., 2022) investigates rate-distortion theory of machine learning models for simple architectures under rather specific assumptions, the information-theoretic limits of neural network compression are generally unknown in practically relevant settings. In this context, the family of compressed models generated by ACIP conveniently provides an empirical (upper) bound on the distortion-rate function of a large-scale model from a single optimization run.
@@ 5 Conclusion
In this work, we have introduced Any Compression via Iterative Pruning (ACIP), a simple end-to-end algorithm to determine the compression-performance trade-off of pre-trained models. The underlying score map ranking allows us to materialize models of any compression rate in real-time. We have demonstrated empirically that the downstream performance of the resulting models is superior to existing, layer-wise factorization approaches. The flexibility and efficiency of ACIP make it a practical tool for deploying large-scale models in resource-constrained settings, especially in combination with other compression techniques such as quantization.
Discussion.
Our main results in Figures˜4 and 5 resemble the well-known phenomenon of scaling laws Kaplan et al. (2020); Hoffmann et al. (2022). Recently, it has been shown that any-size models can be achieved through pre-training Devvrit et al. (2024); Gonzalez & Shivanna (2025) (see also Section˜4), exhibiting similar trade-offs as ACIP. Establishing a rigorous connection between these two fields of research could be a fruitful avenue of future work.
In a similar vein, we observe that more recent models tend to be less compressible (e.g., compare the slopes of LLaMA-13B and Qwen2.5-14B in Figure˜5). We hypothesize that this relates to newer models carrying denser information per weight, since they were trained on much larger datasets Allen-Zhu & Li (2024). Also, the distribution of the calibration dataset (C4 in our case) might play an important role in this context.
A notable technical limitation of our work is that we have only focused on models that are tunable on a single (NVIDIA H100) GPU in bf16-precision. Hence, the scaling behavior of ACIP for larger LLMs (30B+) remains to be explored. We also emphasize that ACIP could be transferred to other modalities, architectures, and tasks without any notable modifications. Finally, a more detailed study of inference speed (beyond the results of Section˜D.1) could provide useful insights into the interplay of low-rank models and their efficiency.
Broader Impact Statement
The primary broader impact of our work lies in increasing the accessibility and practicality of large Foundation Models. By empowering practitioners to effortlessly navigate the compression-performance trade-off without costly recomputation, ACIP helps to democratize the deployment of advanced AI. This could unlock novel applications in resource-scarce domains — from on-device mobile assistants to intelligent systems in manufacturing and automotive sectors — and support researchers with limited computational budgets. Ultimately, our approach contributes to a more sustainable and inclusive AI ecosystem by enabling the efficient use of pre-trained models, reducing both computational and environmental overhead.
In parallel with these benefits, it is crucial to acknowledge the ethical implications and responsibilities that come with democratizing powerful technology. The same accessibility that fosters innovation could also lower the barrier for malicious applications, such as the efficient generation of spam or real-time misinformation on edge devices. This underscores the critical need for the AI/ML community to proactively develop robust safety mechanisms and ethical guidelines that are effective for models of all sizes. By doing so, we can ensure the positive impacts of accessibility are not undermined by potential misuse.
@@ Discussion.
Acknowledgments
@@
We kindly acknowledge funding by the German Federal Ministry of Education and Research (BMBF) within the project “More-with-Less: Effiziente Sprachmodelle für KMUs” (grant no. 01IS23013A). Computational resources were provided by the German AI Service Center WestAI and used to conduct the numerical experiments of this work.
Appendix A Additional Remarks
In this section, we discuss a few additional aspects (in Q&A format) about our method and experimental design that were not (fully) addressed in the main part for the sake of brevity.
@@ Appendix A Additional Remarks
    We argue that these are fundamentally different compression approaches. Full weight manipulations, in principle, have the potential to lead to more powerful compressions because they have more degrees of freedom (analogously to full-weight fine-tuning vs. PEFT). Therefore, they should not be seen as competing methods but complementary ones. We admit that practitioners probably would not favor ACIP over well-established and widely supported quantization techniques. However, the adapter-style nature of ACIP makes it suitable for a combination. This can, for example, allow ACIP to be further improved through a quantization of the singular vector matrices as demonstrated in Section˜3.4.
@@
    Why did you not compare with model distillation or combine ACIP with it?
@@
    While model distillation can lead to outstanding compression results, e.g., see Busbridge et al. (2025); Raschka (2024), this approach requires significantly more resources than ACIP, typically orders of magnitudes more. A direct comparison is therefore not meaningful from our point of view, as it should at least be based on approximately the same computational budget.
@@
    Let us first summarize several benefits of our end-to-end optimization approach from the main paper: (i) it is conceptually simple and requires no feature engineering, (ii) an error correction can be injected with almost no extra costs, (iii) it allows us to perform efficient and accurate Any Compression.
    Apart from that, and to the best of our knowledge, existing compression algorithms that use layer-wise updates like ASVD Yuan et al. (2024), SVD-LLM Wang et al. (2024), or WeLore Jaiswal et al. (2024) require a separate fine-tuning step to achieve competitive downstream performance at stronger compression ratios. Therefore, the lower costs of layer-wise compression are actually dominated by a more expensive backpropagation-based step. It remains open if similar results can be obtained by a fully tuning-free algorithm.
@@
    Committing to a backpropagation-based algorithm (see Q3) means that we have to deal with increased memory requirements. As such, matrix factorization is not helpful in that respect because the number of parameters might even increase initially (for instance, an SVD-parametrization basically doubles the size of a quadratic weight matrix). On the other hand, tuning and pruning only the bottleneck layer (i.e., the singular value masks in case of ACIP) has the potential for drastic size reductions and is highly parameter-efficient. For example, the number of tunable mask parameters for LLaMA-7B with ACIP is <<1M.
    With this in mind, SVD as a specific matrix factorization is an obvious candidate due to its beneficial mathematical and numerical properties, in particular, optimal low-rank matrix approximation and stable matrix operations due to orthogonality.
@@ Appendix B Implementation Details
Following previous work on LLM compression, we use C4 Raffel et al. (2019) for training as it is a good proxy of a general-purpose dataset. In the context of ACIP, it should be primarily seen as a calibration dataset that allows us to propagate meaningful activations through a pre-trained model while performing structured pruning. Overfitting to the distribution of C4 is implicitly mitigated, since we only tune very few parameters (masks and low-rank adapters) compared to the total model size. As loss function $\mathcal{L}$ in (8), we use the standard negative log-likelihood loss for next-token prediction.
@@
All considered (evaluation) datasets and pre-trained models are imported with the HuggingFace transformers-library in bfloat16-precision. Our experiments were implemented with PyTorch Paszke et al. (2019) and the Lightning package.
ACIP-specifics.
@@
As mentioned in Remark˜2.1, we apply a linear scheduler that increases the regularization parameter $\lambda$ dynamically over the pruning process. This ensures that the pruning becomes more and more aggressive over time and the stopping criterion will be reached at some point. Across all experiments, we use $\lambda=1e{-}3$ as initial value and increase it by a factor of $1.01$ every $4$ steps (this amounts to a doubling of $\lambda$ at about every $280$ steps).
@@
As pointed out in Section˜2.2.2, we choose a target compression rate as a stopping criterion for ACIP. In most experiments, a rate of $r_{\text{stop}}=0.4$ is reasonable (i.e., only 40% or the original parameters remain), and we refer to Section˜D.4 for further discussion and analysis. After the stopping criterion is reached, we tune the low-rank adapter for 1k more steps while the masks are frozen (see Section˜2.2.2).
@@
The mask parameters in (7) are rescaled by a fixed factor of $0.02$ to ensure a better alignment with the numerical range of the remaining network weights. The low-rank adapters are created with $r=32$, $\alpha=16$, and dropout $0.05$. For LLaMA-7B, the number of tunable parameters amounts to <<1M mask parameters and approximately 80M low-rank adapter parameters.
@@
For sample data from C4, we use $1024$ tokens per sample and a batch size of $4$. We use Adam Kingma & Ba (2015) as optimizer without weight decay and a learning rate of 5e55e{-}5.
Runtime analysis.
@@
ACIP requires significantly fewer steps than fine-tuning. Depending on when the stopping criterion is reached, it typically takes 1.5k - 2.5k steps, including 1k post-tuning steps of the low-rank adapters. For LLaMA-7B, for example, this amounts to a wall clock runtime of <30<30 minutes, including the initial SVD computations for the base model parametrization. All runs were performed on single NVIDIA H100 GPUs. See also Section˜D.1 for a more detailed efficiency analysis.
Fine-tuning.
@@
In all post-compression fine-tunings (see Section˜3.3), we simply continue training ACIP’s low-rank adapters (the optimizer states are reset). We train for 2525k steps on C4 with a batch size of $4$ and a learning rate of 2e42e{-}4.
Appendix C Supplementary Results for Section˜3.1Section˜3.4
@@
Figure˜A8 complements the trade-off curves in Figures˜4 and 5 by all other considered evaluation metrics (see Section˜3.1). Table˜A2 reports these results in terms of numbers, including all fine-tuning results for all models (see Section˜3.3). Table˜A3 analyzes the effect of fine-tuning of ACIP compared to existing SVD-based compression methods. Table˜A4 provides more detailed evaluation results on fine-tuning a quantized and compressed LLaMA-7B model (see Section˜3.4).
(a) WikiText-2
(b) ARC Challenge
(c) ARC Easy
(d) HellaSwag
(e) MathQA
(f) OpenBookQA
(g) PIQA
(h) Winogrande
Figure A8: Compression-performance trade-off curves generated by ACIP on WikiText-2 and individual LM-Eval tasks, complementing the results of Figures˜4 and 5.
@@
Table A2: Evaluation results for ACIP on all considered LLMs. Scores on C4 and WikiText-2 are measured in perplexity (smaller is better), and the LM-Eval zero-shot tasks are measured in accuracy (higher is better). The results of LLaMA2-13B were achieved by ignoring all up-projection layers in ACIP (see Section˜D.8).
@@
Orig. 9.99 5.05 0.56 0.82 0.63 0.53 0.35 0.81 0.75 0.64
Table A3: Evaluation of LLaMA-7B on WikiText-2 (perplexity, smaller is better) under different compression ratios, with and without post-training fine-tuning. We compare ACIP with the existing SVD-based compression methods ASVD Yuan et al. (2024) and SVD-LLM Wang et al. (2024), see also Section˜3.2. The scores for ASVD and SVD-LLM are taken from Wang et al. (2024, Table 4). Note that ACIP was fine-tuned on C4, while ASVD and SVD-LLM fine-tuned on WikiText-2 directly.
Compression Ratio 40% 50% 60% 70% 80%
Method
ASVD 57057.00 15358.00 01407.00 00051.00 00011.14
ASVD + LoRA FT 44.81 21.83 14.86 10.16 8.37
SVD-LLM 42.30 23.97 13.11 9.56 7.94
SVD-LLM + LoRA FT 17.93 13.26 10.65 9.14 7.78
ACIP 24.00 16.17 12.46 10.34 8.83
ACIP + FT 17.33 11.88 9.63 8.27 7.19
Table A4: More detailed evaluation results for our quantization experiments in Section˜3.4, reported in terms of numbers.
Eff. model size [GB] C4 WikiText-2 ARC_c ARC_e HellaS. MathQA Openb. PIQA WinoG. LM Eval
Size Ablation Avg. ↑